OpenStudy (anonymous):

A student increases the temperature of a 417 cm3 balloon from 278 K to 308 K. Assuming constant pressure, what should the new volume of the balloon be? (Points : 5)

OpenStudy (anonymous):

Use PV=nRT Assuming the P and the n are constant V/T=nR/P You know the left part won't change, so you have to solve the equation V/T=V'/T'

OpenStudy (anonymous):

417 cm3 462 cm3 924 cm3 376 cm3 this is my answers and i chose the last one am i correct?

OpenStudy (anonymous):

How much heat do you need to raise the temperature of 100 g of aluminum from 30°C to 150°C? (Points : 5) 10.8 kJ 10.8 × 103 kJ 10.8 kJ 10.8 × 103 kJ

OpenStudy (anonymous):

my answer is 10.8kj

OpenStudy (anonymous):

For the first question you're wrong, 417/278=462/308, so 462 is the right answer